# LoCo: Low-Contrast-Enhanced Contrastive Learning for Semi-Supervised Endoscopic Image Segmentation

### v1

@@ Abstract
1 Introduction
[…]
Electronic endoscope is indispensable for diagnosing tumors in the throat and digestive tract, offering exceptionally high clarity, resolution, and vivid color imaging to detect early disease signs—especially the small, concealed lesions often associated with malignant tumors [1, 2]. However, current tumor detection and diagnostic accuracy are constrained by the physician’s expertise and experience level. Precisely segmenting tumor lesions can yield valuable annotated information on affected areas, as well as refined analyses of lesion size, shape, and color [3]. This capability is essential for accurately identifying and diagnosing tumor abnormalities, enabling more effective tumor screening, personalized treatment planning, and ultimately enhancing patient outcomes.
@@ 1 Introduction
Recently, remarkable advancements in deep learning models have led to notable improvements in semantic segmentation across diverse domains, yielding promising outcomes [4, 5, 6, 7, 8, 9, 10]. However, these methods generally rely on extensive manual annotations to attain competitive performance. In endoscopic applications, the limitation of the annotations constrains model performance on endoscopic image datasets. Semi-supervised learning offers an efficient solution by leveraging limited labeled data with abundant unlabeled data for training. Current segmentation methods [11, 12, 13, 14] primarily integrate consistency learning [15] with pseudo-labeling [16]. These methods utilize model predictions on weakly perturbed unlabeled data as pseudo-labels to supervise predictions on strongly perturbed data, while applying a confidence threshold to filter out low-confidence pseudo-labels, thus reducing confirmation bias [17]. Additionally, segmentation foundation models, such as MedSAM [18], offer an effective approach to mitigating the challenge of limited annotations. This is due to their strong generalization capabilities, which allow them to perform well on downstream tasks with only a small amount of data for fine-tuning. Despite these advances, the issue of low contrast in endoscopic images still poses a significant challenge for the model to accurately segment regions corresponding to normal, benign, and malignant tissue.
@@
The low contrast issue can be divided into two categories: low boundary contrast and low inter-class contrast (Fig. 1). Low boundary contrast refers to the low contrast caused by the high similarity between normal and tumor tissues, along with blurred boundaries of tumors. Low inter-class contrast refers to the low contrast resulting from the high similarity between different types of tumors. Regarding the low boundary contrast, several studies have integrated boundary information to bolster the model’s ability to differentiate between normal and tumor regions. Notable methods include PraNet [19], BCNet [20], and ACL-Net [21]. For instance, Polyper [22] introduced boundary-sensitive attention mechanisms to enhance segmentation accuracy by refining features near boundary regions. Regarding the low inter-class contrast, there is little research addressing this issue. This is primarily because most existing endoscopic datasets contain only two categories: background and tumor, without further annotation of different tumor types. As a result, related studies are unable to effectively propose a solution for low inter-class contrast. To bridge this gap, this work utilizes a large proprietary dataset collected from the First Affiliated Hospital, Sun Yat-Sen University. This dataset includes pixel-level annotations for three categories: normal tissue, benign tumor, and malignant tumor, facilitating research aimed at addressing the challenge of low inter-class contrast.
@@
In this study, we propose a novel semi-supervised segmentation framework, termed LoCo, which employs low-contrast-enhanced contrastive learning (LCC) for endoscopic image segmentation. Specifically, to facilitate the model’s learning of discriminative features, we propose two strategies: inter-class contrast enhancement (ICE) and boundary contrast enhancement (BCE) for selecting low-contrast pixels to construct positive and negative pairs for contrastive learning. In the ICE, we identify low-contrast pixels that exhibit the lowest similarity to class embeddings within each category for contrastive learning, enabling the model to establish a more robust decision boundary for effectively differentiating between various tumor types. In the BCE, we identify low-contrast boundary pixels with the highest boundary feature similarity for contrastive learning, which enhances the distinct separation of tumor boundaries from normal tissue. Furthermore, to tackle the challenges of low utilization of pseudo-labels, we design a confidence-based dynamic filter. This filter dynamically adjusts thresholds for different categories based on the class-wise confidence of pseudo-labels, thereby enhancing the utilization of pseudo-labels for minority classes. Experimental results from a proprietary laryngeal cancer dataset and two polyp segmentation datasets indicate that our proposed LoCo framework achieves state-of-the-art performance, exceeding existing methods by a significant margin. The primary contributions of our work can be summarized as follows:
@@
The remainder of the paper is organized as follows. Section 2 provides a brief review of related work pertinent to this article. Section 3 details the proposed LoCo method. Section 4 presents and analyzes the relevant experimental results. Finally, Section 5 reviews and summarizes the article.
2 Related Works
2.1 Semi-supervised semantic segmentation
@@
Semi-supervised semantic segmentation aims at harnessing the rich visual information in unlabeled images to improve segmentation accuracy. Recently, this approach has achieved notable success in medical image analysis [23, 24, 25, 26, 27, 28] and remote sensing interpretation [29, 30, 31]. Most studies are inspired by advancements in semi-supervised learning, proposing straightforward yet effective designs centered around entropy minimization [32, 33, 34] and consistency regularization [15, 35, 36]. Entropy minimization aims to assign pseudo-labels to unlabeled data using knowledge from provided labels, followed by re-training the model on the combined dataset, as seen in methods like CPS [37], CCVC [38], and St++ [39]. Despite its effectiveness, the offline self-training pipeline is infrequently employed due to the complexities of its three-stage training process. Consequently, many recent studies favor end-to-end pseudo-labeling frameworks powered by consistency regularization, which emphasizes the design of perturbations on input data and mandates that the model produces consistent predictions for the same data across different augmentations. Notable examples include MixMatch [35], AugSeg [12], and UniMatch [13]. However, these methods neglect the issue of low contrast, leading to suboptimal segmentation performance on endoscopic images.
2.2 Contrastive learning
@@
Contrastive learning is an effective method for representation learning. It encourages the model to learn better representations by minimizing the distance of positive pairs while maximizing the distance of negative pairs. This method has demonstrated impressive performance in both self-supervised [40, 41, 42, 43, 44] and supervised [45] learning. Recently, contrastive learning has also been introduced to segmentation tasks. Most of those methods, such as P2Seg [46], C3-SemiSeg [47], U2PL [48], and PLGCL [25], leverage contrastive learning to mine more effective supervision, thereby enabling the model to learn more discriminative features. Despite the improvement in segmentation performance, some studies [11, 49] suggest that selecting more appropriate samples (e.g., samples at the edge of the distribution) to construct positive and negative pairs plays a crucial role in learning clearer decision boundaries. In our proposed low-contrast-enhanced contrastive learning, we select low-contrast pixels from both the category and boundary perspectives to construct positive and negative pairs. Then, we leverage the proposed low-contrast enhanced contrastive loss to optimize the distance of those pairs, which allows the model to learn more contrastive features in endoscopic images.
@@
3 Methods
3.1 Overall LoCo
Fig.2 illustrates the overall architecture of LoCo, which is based on the mean-teacher framework. The LoCo comprises three primary components: a student network, a teacher network, and a Low-contrast-enhanced contrastive learning module (i.e., LCC, see Section 3.2). For the teacher network, it utilizes the same architecture as the student network, updating its parameters via an Exponential Moving Average (EMA) scheme. The EMA is defined as follows:
@@ 3.1 Overall LoCo
where $\theta_{s}$ and $\theta_{t}$ represent the learned parameters of the student network and the teacher network, respectively. $\alpha\in[0,1]$ is a momentum hyperparameter, and it is typically set to 0.999 by default [50, 12]. Regarding the student network, its parameters $\theta_{s}$ are updated with the optimization of Stochastic Gradient Descent (SGD). In the medical domain, the availability of labeled medical images is often limited, while the volume of unlabeled images can be substantial. This scarcity of labeled data hampers the model’s ability to achieve satisfactory segmentation results. Consequently, effectively utilizing the abundant unlabeled data is crucial for training a robust model. In this study, we utilize the teacher network to generate reliable pseudo-labels through a Confidence-based Dynamic Filter (i.e., CDF, see Section 3.3) for unlabeled data. These pseudo-labels, along with labeled images, are combined to supervise the learning of the student network. Additionally, LCC is employed to improve segmentation performance in low-contrast images through inter-class contrast enhancement (i.e., ICE, see Section 3.2.1) and boundary contrast enhancement (i.e., BCE, see Section 3.2.2).
@@
3.2 LCC
The goal of LCC is to mine low-contrast pixels and enhance the model’s ability to distinguish these pixels via contrastive learning. Specifically, We first select low-contrast pixels from the perspectives of both class (i.e., tumor types) and tumor boundary. Each selected low-contrast pixel $x_{i,j}$’s corresponding feature embedding $z_{i,j}$ will be added to the low-contrast feature embedding set $\mathcal{H}$ (detail refer to section 3.2.1 and 3.2.2). Then, given a feature embedding $z_{i,j}\in\mathcal{H}$ belongs to class $c$, we guide the model to learn discriminative features by pulling $z_{i,j}$ close to the class embedding $\bar{z}^{c}$, while pushing it further away from the other class embeddings not belong to class $c$. Formally, let $\{z^{c}\}=\{z_{i,j}^{l}|y_{i,j}^{l}=c,z_{i,j}^{l}\in\mathcal{H}\}\cup\{z_{i,j}^{u}|y_{i,j}^{p}=c,z_{i,j}^{u}\in\mathcal{H}\}$ denote the set of low-contrast feature embeddings belong to class $c$. The loss of LCC can be defined as:
@@ 3.2.2 BCE
3.3 CDF
In semi-supervised segmentation, it is essential to use a threshold to filter out low-confidence pseudo-labels and prevent confirmation bias. A common practice is to employ a fixed threshold. However, this approach has two significant drawbacks: 1) It can result in the loss of many valuable pseudo-labels during the early stages of training. 2) In endoscopic image segmentation, the issue of pixel imbalance is severe. Certain minority classes, such as benign tumor, account for only a small portion of the total pixels in all images. This often results in lower confidence for these minority classes, ultimately leading to extremely low utilization of pseudo-labels for minority classes. (see section 4.3).
[…]
@@ 3.3 CDF
Malignant Benign mIoU Malignant Benign mIoU Malignant Benign mIoU
Supervised 65.69 36.65 51.17 70.98 45.96 58.47 73.47 55.63 64.55
FixMatch 70.20 46.44 58.32 73.93 49.00 61.46 73.62 54.38 64.00
U2PL 68.33 46.68 57.50 72.31 50.81 61.56 73.93 57.98 65.95
UniMatch 70.08 47.88 58.98 73.67 49.11 61.39 74.42 55.13 64.77
AugSeg 71.21 48.83 60.02 74.22 51.22 62.72 74.53 57.76 66.14
CorrMatch 67.09 41.92 54.50 71.11 44.37 57.74 71.21 48.14 59.67
Ours 72.19 52.96 62.57 75.75 56.51 66.13 76.23 61.56 68.89
@@
4 Experiments
4.1 Experiment setup
4.1.1 Dataset
@@
4.1.2 Evaluation
To evaluate the performance of our method in the semi-supervised setting, we conduct experiments under $10\%$, $30\%$, and $50\%$ partition protocols (i.e., only $10\%$, $30\%$, and $50\%$ of labeled data are available). We adopt the mean of intersection of union (mIoU) as our evaluation metric. Additionally, on FAHSYU-LC, to evaluate the model’s segmentation performance on different tumor types, we also report the class intersection of union (IoU) for each category. On Kvasir-SEG and CVC-ClinicDB, we only report the mIoU since there are only two categories (i.e., background and tumor).
4.1.3 Implementation details
@@
4.2 Comparison with state-of-the-art methods
@@
Table 1 reports the experiment results on FAHSYU-LC. The results show that our method consistently outperforms other state-of-the-art methods. Under $10\%$, $30\%$, and $50\%$ label partitions, our method outperforms the previous best by $2.55\%$, $3.41\%$, and $2.75\%$. Moreover, our method brings greater improvements on minority classes such as benign, which surpasses the previous best approach by $4.13\%$, $5.29\%$, and $3.58\%$ on each split.
Table 2 shows the comparison results with the state-of-the-art methods on Kvasir-SEG. It is clear that our method consistently outperforms the second-best method by $2.57\%$, $1.61\%$, and $1.83\%$ on Kvasir-SEG. Table 3 shows the comparison results on CVC-ClinicDB., our method archives $83.81\%$, $88.49\%$, and $92.05\%$ under $10\%$, $30\%$, and $50\%$ partition protocols, which is on par with the state-of-the-art methods.
Table 2: Segmentation Results on Kvasir-SEG. Bold denotes the best performance, while underline denotes the second.
[…]
      10%       30%       50%
      Supervised       74.74       83.30       83.79
      FixMatch       83.05       84.57       84.60
      U2PL       82.38       84.36       85.40
      UniMatch       83.16       82.38       84.42
      AugSeg       83.61       85.16       85.27
      CorrMatch       81.18       84.38       85.08
      Ours       86.18       86.77       87.23
Table 3: Segmentation results on CVC-ClinicDB. Bold denotes the best performance, while underline denotes the second.
      Method       mIoU (%)
      10%       30%       50%
      Supervised       75.09       82.66       88.51
      FixMatch       82.63       86.08       84.68
      U2PL       83.51       86.92       87.64
      UniMatch       82.66       86.00       84.87
      AugSeg       84.13       86.95       89.57
      CorrMatch       80.47       83.60       83.47
      Our       83.81       88.49       92.05
4.3 Ablation Studies
Table 4: Ablation results of LoCo on FAHSYU-LC under 10% label partition.
[…]
M1 65.69 36.65 51.17
M2 70.42 45.68 58.05
M3 69.91 47.49 58.70
M4 70.18 49.04 59.61
M5 71.85 51.43 61.64
M6 71.99 51.98 61.98
M7 72.19 52.96 62.57
Table 4 presents the ablation results of the LoCo method on the FAHSYU-LC dataset under a 10% label partition. It compares the performance of seven distinct model variants, labeled M1 through M7, as detailed below:

### v2

# Low-Contrast-Enhanced Contrastive Learning for Semi-Supervised Endoscopic Image Segmentation

@@ Abstract
I Introduction
Electronic endoscope is indispensable for diagnosing tumors in the throat and digestive tract, offering exceptionally high clarity, resolution, and vivid color imaging to detect early disease signs—especially the small, concealed lesions often associated with malignant tumors [1]. However, current tumor detection and diagnostic accuracy are constrained by the physician’s expertise and experience level. Precisely segmenting tumor lesions can yield valuable annotated information on affected areas, as well as refined analyses of lesion size, shape, and color [2]. This capability is essential for accurately identifying and diagnosing tumor abnormalities, enabling more effective tumor screening, personalized treatment planning, and ultimately enhancing patient outcomes.
@@ I Introduction
Recently, remarkable advancements in deep learning models have led to notable improvements in semantic segmentation across diverse domains, yielding promising outcomes [3, 4, 5, 6, 7, 8]. However, these methods generally rely on extensive manual annotations to attain competitive performance. In endoscopic applications, the limitation of the annotations constrains model performance on endoscopic image datasets. Semi-supervised learning offers an efficient solution by leveraging limited labeled data with abundant unlabeled data for training. Current segmentation methods [9, 10, 11, 12] primarily integrate consistency learning [13] with pseudo-labeling [14]. These methods utilize model predictions on weakly perturbed unlabeled data as pseudo-labels to supervise predictions on strongly perturbed data, while applying a confidence threshold to filter out low-confidence pseudo-labels, thus reducing confirmation bias [15]. Additionally, segmentation foundation models, such as MedSAM [16], offer an effective approach to mitigating the challenge of limited annotations. This is due to their strong generalization capabilities, which allow them to perform well on downstream tasks with only a small amount of data for fine-tuning. Despite these advances, the issue of low contrast in endoscopic images still poses a significant challenge for the model to accurately segment regions corresponding to normal, benign, and malignant tissues.
@@
The issue of low contrast can be classified into two categories: low boundary contrast and low inter-class contrast (see Fig. 1). Low boundary contrast arises from the high similarity between normal and tumor tissues, compounded by the blurred boundaries of tumors. In contrast, low inter-class contrast is characterized by the high similarity among different tumor types. To address low boundary contrast, several studies have incorporated boundary information to enhance the model’s capacity to differentiate between normal and tumor regions. Notable approaches include PraNet [17], BCNet [18], and ACL-Net [19]. For example, Polyper [20] introduced boundary-sensitive attention mechanisms, improving segmentation accuracy by refining features in proximity to boundary regions. Conversely, there is limited research focused on low inter-class contrast. This scarcity is largely due to the fact that most existing endoscopic datasets feature only two categories: background and tumor, lacking detailed annotations for various tumor types. Consequently, related studies face challenges in effectively addressing the issue of low inter-class contrast. To bridge the gap, this work utilizes a large proprietary dataset collected from the First Affiliated Hospital, Sun Yat-sen University. This dataset includes pixel-level annotations for three categories: normal tissue, benign tumors, and malignant tumors, thereby facilitating research aimed at tackling the challenge of low inter-class contrast.
@@
In this study, we propose a novel semi-supervised segmentation framework, termed LoCo, which employs low-contrast-enhanced contrastive learning (LCC) for endoscopic image segmentation. Specifically, to facilitate the model’s learning of discriminative features, we propose two strategies: inter-class contrast enhancement (ICE) and boundary contrast enhancement (BCE) for selecting low-contrast pixels to construct positive and negative pairs for contrastive learning. In the ICE, we identify low-contrast pixels that exhibit the lowest similarity to class embeddings within each category for contrastive learning, enabling the model to establish a more robust decision boundary for effectively differentiating between various tumor types. In the BCE, we identify low-contrast boundary pixels with the highest boundary feature similarity for contrastive learning, which enhances the distinct separation of tumor boundaries from normal tissue. Furthermore, to tackle the challenges of low utilization of pseudo-labels, we design a confidence-based dynamic filter (CDF). This filter dynamically adjusts thresholds for different categories based on the class-wise confidence of pseudo-labels, thereby enhancing the utilization of pseudo-labels for minority classes. Experimental results from a proprietary laryngeal cancer dataset and two polyp segmentation datasets indicate that our proposed LoCo framework achieves state-of-the-art performance, exceeding existing methods by a significant margin. The primary contributions of our work can be summarized as follows:
@@
The remainder of the paper is organized as follows. Section II provides a brief review of related work pertinent to this article. Section III details the proposed LoCo method. Section IV presents and analyzes the relevant experimental results. Finally, Section V reviews and summarizes the article.
II Related Works
II-A Semi-supervised semantic segmentation
@@
Semi-supervised semantic segmentation seeks to leverage the abundant visual information found in unlabeled images to enhance segmentation accuracy [21]. Recently, these approaches have made significant strides in fields such as medical image analysis [22, 23, 24, 25, 26] and remote sensing interpretation [27, 28, 29]. Most studies draw inspiration from advancements in semi-supervised learning, proposing straightforward yet effective designs centered around entropy minimization [30, 31, 32] and consistency regularization [13, 33]. Entropy minimization involves assigning pseudo-labels to unlabeled data by utilizing the knowledge derived from labeled samples, followed by retraining the model on the combined dataset. This approach is exemplified by methods such as CPS [34], CCVC [35], and St++ [36]. Despite its effectiveness, the offline self-training pipeline is rarely implemented due to the complexities associated with its three-stage training process. As a result, many studies prefer end-to-end pseudo-labeling frameworks driven by consistency regularization, which focus on generating perturbations in input data and require the model to produce consistent predictions for the same data across various augmentations. Notable examples include AugSeg [10], UniMatch [11], and DDFP [37]. However, these methods often overlook the issue of low contrast, which can lead to suboptimal segmentation performance, particularly in endoscopic images.
II-B Contrastive learning
@@
Contrastive learning is a powerful approach for representation learning, designed to enhance the model’s ability to generate robust representations by minimizing the distance between positive pairs while maximizing the distance between negative pairs. This technique has shown remarkable success in both self-supervised [38, 39, 40, 41, 42] and supervised [43] learning contexts. Recently, contrastive learning has been introduced to segmentation tasks, with methods such as P2Seg [44], C3-SemiSeg [45], U2PL [46], and PLGCL [24] utilizing this framework to extract more effective supervisory signals, thereby enabling the model to learn more discriminative features. While these advancements have led to improved segmentation performance, some studies [9, 47] indicate that the careful selection of samples—particularly those at the edges of the distribution—for constructing positive and negative pairs is critical for learning clearer decision boundaries. In our proposed low-contrast-enhanced contrastive learning approach, we specifically select low-contrast pixels from both categorical and boundary perspectives to form positive and negative pairs. We then employ a novel low-contrast-enhanced contrastive loss to optimize the distances of these pairs, facilitating the model’s ability to learn more pronounced contrastive features in endoscopic images.
@@
III Methods
III-A Overall LoCo
Fig.2 illustrates the overall architecture of LoCo, which is based on the mean-teacher framework. The LoCo comprises three primary components: a student network, a teacher network, and a Low-contrast-enhanced contrastive learning module (i.e., LCC, see Section III-B). For the teacher network, it utilizes the same architecture as the student network, updating its parameters via an Exponential Moving Average (EMA) scheme. The EMA is defined as follows:
@@ III-A Overall LoCo
where $\theta_{s}$ and $\theta_{t}$ represent the learned parameters of the student network and the teacher network, respectively. $\alpha\in[0,1]$ is a momentum hyperparameter, and it is typically set to 0.999 by default [48, 10]. Regarding the student network, its parameters $\theta_{s}$ are updated with the optimization of Stochastic Gradient Descent (SGD). In the medical domain, the availability of labeled medical images is often limited, while the volume of unlabeled images can be substantial. This scarcity of labeled data hampers the model’s ability to achieve satisfactory segmentation results. Consequently, effectively utilizing the abundant unlabeled data is crucial for training a robust model. In this study, we utilize the teacher network to generate reliable pseudo-labels through a Confidence-based Dynamic Filter (i.e., CDF, see Section III-C) for unlabeled data. These pseudo-labels, along with labeled images, are combined to supervise the learning of the student network. Additionally, LCC is employed to improve segmentation performance in low-contrast images through inter-class contrast enhancement (i.e., ICE, see Section III-B1) and boundary contrast enhancement (i.e., BCE, see Section III-B2).
@@
III-B LCC
The goal of LCC is to mine low-contrast pixels and enhance the model’s ability to distinguish these pixels via contrastive learning. Specifically, We first select low-contrast pixels from the perspectives of both class (i.e., tumor types) and tumor boundary. Each selected low-contrast pixel $x_{i,j}$’s corresponding feature embedding $z_{i,j}$ will be added to the low-contrast feature embedding set $\mathcal{H}$ (detail refer to section III-B1 and III-B2). Then, given a feature embedding $z_{i,j}\in\mathcal{H}$ belongs to class $c$, we guide the model to learn discriminative features by pulling $z_{i,j}$ close to the class embedding $\bar{z}^{c}$, while pushing it further away from the other class embeddings not belong to class $c$. Formally, let $\{z^{c}\}=\{z_{i,j}^{l}|y_{i,j}^{l}=c,z_{i,j}^{l}\in\mathcal{H}\}\cup\{z_{i,j}^{u}|y_{i,j}^{p}=c,z_{i,j}^{u}\in\mathcal{H}\}$ denote the set of low-contrast feature embeddings belong to class $c$. The loss of LCC can be defined as:
@@ III-B2 BCE
III-C CDF
In semi-supervised segmentation, it is essential to use a threshold to filter out low-confidence pseudo-labels and prevent confirmation bias. A common practice is to employ a fixed threshold. However, this approach has two significant drawbacks: 1) It can result in the loss of many valuable pseudo-labels during the early stages of training. 2) In endoscopic image segmentation, the issue of pixel imbalance is severe. Certain minority classes, such as benign tumor, account for only a small portion of the total pixels in all images. This often results in lower confidence for these minority classes, ultimately leading to extremely low utilization of pseudo-labels for minority classes. (see section IV-C).
TABLE I: Segmentation results on FAHSYU-LC. Bold denotes the best performance, while underline denotes the second.
@@ III-C CDF
IoU(M) IoU(B) mIoU DSC NSD IoU(M) IoU(B) mIoU DSC NSD IoU(M) IoU(B) mIoU DSC NSD
Supervised 65.69 36.65 51.17 58.17 60.38 70.98 45.96 58.47 67.29 69.78 73.47 55.63 64.55 71.02 73.60
FixMatch 70.20 46.44 58.32 66.79 69.21 73.93 49.00 61.46 70.18 72.71 73.62 54.38 64.00 70.08 72.67
U2PL 68.33 46.68 57.50 65.13 67.54 72.31 50.81 61.56 71.02 73.66 73.93 57.98 65.95 72.56 75.26
UniMatch 70.08 47.88 58.98 66.77 69.28 73.67 49.11 61.39 69.84 72.36 74.42 55.13 64.77 71.06 73.61
AugSeg 71.21 48.83 60.02 68.06 70.56 74.22 51.22 62.72 71.80 74.42 74.53 57.76 66.14 73.20 75.89
CorrMatch 67.09 41.92 54.50 64.77 67.08 71.11 44.37 57.74 68.34 70.80 71.21 48.14 59.67 67.63 70.07
Ours 72.19 52.96 62.57 70.49 73.13 75.75 56.51 66.13 75.54 78.38 76.23 61.56 68.89 75.82 78.58
@@
IV Experiments
IV-A Experiment setup
IV-A1 Dataset
@@
IV-A2 Evaluation
To evaluate the performance of our method in the semi-supervised setting, we conduct experiments under $10\%$, $30\%$, and $50\%$ partition protocols (i.e., only $10\%$, $30\%$, and $50\%$ of labeled data are available). We adopt the mean of intersection of union (mIoU), dice similarity coefficient (DSC), and normalized surface distance (NSD) as our evaluation metric. Additionally, on FAHSYU-LC, to evaluate the model’s segmentation performance on different tumor types, we also report the class intersection of union (IoU) for each category. Here, ”IoU(M)” represents the IoU for malignant tumor and ”IoU(B)” represents the IoU for benign tumor. On Kvasir-SEG and CVC-ClinicDB, we only report the mIoU, DSC, and NSD since there are only two categories (i.e., background and tumor).
IV-A3 Implementation details
@@
IV-B Comparison with state-of-the-art methods
@@
Table I reports the experiment results on FAHSYU-LC. The results show that our method consistently outperforms other state-of-the-art methods. Under $10\%$, $30\%$, and $50\%$ label partitions, our method outperforms the previous best by $2.55\%$, $3.41\%$, and $2.75\%$ in terms of mIoU. Moreover, our method brings greater improvements on minority classes such as benign, which surpasses the previous best approach by $4.13\%$, $5.29\%$, and $3.58\%$ on each split.
Table II shows the comparison results with the state-of-the-art methods on Kvasir-SEG. It is clear that our method consistently outperforms the second-best method by $2.57\%$, $1.61\%$, and $1.83\%$ on Kvasir-SEG. Table III shows the comparison results on CVC-ClinicDB, our method archives the mIoU of $83.81\%$, $88.49\%$, and $92.05\%$ under $10\%$, $30\%$, and $50\%$ partition protocols, which is on par with the state-of-the-art methods.
TABLE II: Segmentation results on Kvasir-SEG. Bold denotes the best performance, while underline denotes the second.
Method 10% 30% 50%
mIoU DSC NSD mIoU DSC NSD mIoU DSC NSD
Supervised 74.74 76.69 78.83 83.38 82.66 84.73 83.79 83.82 85.93
FixMatch 83.05 82.38 84.44 84.57 83.99 86.02 84.60 84.05 86.14
U2PL 82.38 81.89 83.83 84.36 84.35 86.44 85.40 85.35 87.40
UniMatch 83.16 83.33 85.41 82.37 83.86 86.00 84.41 84.53 86.55
AugSeg 83.61 84.39 86.42 85.16 84.44 86.53 85.27 84.50 86.44
CorrMatch 81.18 81.77 83.83 84.38 83.97 86.18 85.14 83.67 85.76
Ours 86.18 84.92 86.85 86.77 86.03 87.99 87.23 86.18 88.08
TABLE III: Segmentation results on CVC-ClinicDB. Bold denotes the best performance, while underline denotes the second.
Method 10% 30% 50%
mIoU DSC NSD mIoU DSC NSD mIoU DSC NSD
Supervised 76.23 61.24 65.20 82.01 72.49 76.43 88.51 81.15 85.01
FixMatch 82.63 75.35 79.06 86.08 78.32 81.85 84.68 76.80 80.42
U2PL 83.51 76.58 80.03 86.92 80.88 84.43 87.64 80.77 84.58
UniMatch 82.66 74.11 77.66 86.00 78.34 81.89 84.87 75.43 79.47
AugSeg 84.13 78.49 81.81 86.95 81.00 84.89 89.57 80.12 83.45
CorrMatch 80.47 71.78 75.88 83.60 76.28 80.35 83.63 75.10 79.16
Ours 83.81 76.31 79.67 88.48 83.73 87.29 92.05 85.16 88.72
IV-C Ablation Studies
TABLE IV: Ablation results of LoCo on FAHSYU-LC under 10% label partition.
Variant $\mathcal{L}_{sup}$ $\mathcal{L}_{u}$ CDF BCE ICE IoU(M) IoU(B) mIOU DSC NSD
M1 65.69 36.65 51.17 58.17 60.38
M2 70.42 45.68 58.05 66.76 69.28
M3 70.18 49.04 59.61 67.46 69.99
M4 69.91 47.49 58.70 67.35 69.84
M5 71.85 51.43 61.64 70.00 72.61
M6 72.04 52.35 62.20 70.47 73.06
M7 72.19 52.96 62.57 70.49 73.13
Table IV presents the ablation results of the LoCo method on the FAHSYU-LC dataset under a 10% label partition. It compares the performance of seven distinct model variants, labeled M1 through M7, as detailed below: